### v1

@@ Abstract
Library migration is the process of replacing a library with a similar one in a software project. Manual library migration is time consuming and error prone, as it requires developers to understand the Application Programming Interfaces (API) of both libraries, map equivalent APIs, and perform the necessary code transformations. Due to the difficulty of the library migration process, most of the existing automated techniques and tooling stop at the API mapping stage or support a limited set of libraries and code transformations. In this paper, we develop an end-to-end solution that can automatically migrate code between any arbitrary pair of Python libraries that provide similar functionality. Due to the promising capabilities of Large Language Models (LLMs) in code generation and transformation, we use LLMs as the primary engine for migration. Before building the tool, we first study the capabilities of LLMs for library migration on a benchmark of 321 real-world library migrations. We find that LLMs can effectively perform library migration, but some post-processing steps can further improve the performance. Based on this, we develop PyMigTool, a command line application that combines the power of LLMs, static analysis, and dynamic analysis to provide accurate library migration. We evaluate PyMigTool on 717 real-world Python applications that are not from our benchmark. We find that PyMigTool can migrate 32% of the migrations with complete correctness. Of the remaining migrations, only 14% of the migration-related changes are left for developers to fix for more than half of the projects.
@@ 1 Introduction
Library migration is the process of replacing a library with an alternative one in a software project. The process involves identifying the library usage in the codebase, finding the corresponding Application Programming Interfaces (APIs) in the new library, and replacing the old API usage with the new API usage. This can be tedious and error-prone, especially when the library is extensively used in a large codebase kula2018developers .
@@
Researchers have proposed various techniques to automate the library migration processes. Most of these techniques, however, are limited to only finding the analogous APIs (API mapping) teyton2013automatic ; alrubaye2018automating ; alrubaye2019use ; miningAnalogicalAPIs ; zhang2020deep ; di2025deepmig , but do not replace the old API usage in the codebase with the new one (transformation). To find analogous APIs, some of these techniques mine existing migrations teyton2013automatic ; alrubaye2018automating ; alrubaye2019use while others use machine learning models miningAnalogicalAPIs ; di2025deepmig and natural language processing techniques ni2021soar . To the best of our knowledge, there are two techniques that perform transformation ni2021soar ; nikolov2025google but both techniques evaluate their approaches on a small set of library pairs without providing a tool that can be used in practice.
@@
In recent years, Large Language Models (LLMs) have shown that they can be effective in various software engineering tasks ozkaya2023application ; fan2023large ; wang2024software , including the necessary upstream tasks for library migration, such as code generation nguyen2022empirical ; peng2023impact ; murali2023codecompose , code comprehension nam2024using , and code transformation wei2023copiloting ; xia2023automated ; zhou2023hybrid . This suggests that LLMs may be capable of performing library migration.
In this paper, we leverage LLMs to explore and address the challenges of automating library migration in two phases. First, we conduct an experiment using an existing library migration benchmark, PyMigBench pymigbench , to determine the capability and challenges of using LLMs for library migration. We find that LLMs are generally capable of performing migration; however, some post processing can further improve the migration results. Second, based on our findings in first phase, we build a CLI tool, PyMigTool, that automates the full library migration process by combining the power of LLMs with post processing using static and dynamic program analysis. We run an extensive evaluation of PyMigTool on a set of 717 new, simulated migrations.
Phase one – Explore LLMs:
In this phase, we investigate the capabilities and challenges of using LLMs for library migration. We use LLama 3.1, GPT-4o mini, and GPT-4o to perform 321 migrations from PyMigBench. We then measure the correctness of the migrations by comparing the LLM’s migrations with the original developer’s migrations. We find that the three LLMs can correctly migrate 26%, 49%, and 57% of the migrations respectively. Overall, we find that LLMs are capable of library migration, and as expected, the more powerful LLM (GPT-4o) performs better than the less powerful LLMs (LLama 3.1 and GPT-4o mini). We also identify two major challenges that make solely relying on LLMs in a practical tool ineffecive: (1) We provide the LLM with only code files that directly use the source library. However, for a given project, there may be other files that depend on the migrated code and also require changes during the migration. (2) All LLMs often return only part of the code, skipping parts of the code that do not require migration. These limitations, however, can be addressed by post-processing the LLM’s results. We use these insights to improve the migration process in the next phase.
Phase two – Build PyMigTool:
@@
In this phase, we build PyMigTool, a Command Line Interface (CLI) that automates the full library migration process using LLMs and post processing using static and dynamic program analysis. PyMigTool takes as 5 inputs from the user (1) A project that requires migration, with an existing test suite with high coverage. (2) A requirements file that lists the dependencies of the project. (3) A source library that is used in the project. (4) A target library and its version that the user wants to migrate to. (5) Name of an LLM that will be used for the migration. By building a dynamic call graph, PyMigTool identifies all files that require migration and sends those to the LLM. PyMigTool then verifies the results of the migration by comparing the test results before and after migration. If any previously passing tests now fail, PyMigTool performs a Post-processing round where it checks if the LLM skipped any code and merges it back. It also detects if any of functions were made async during the migration, and if so, it finds other parts of the project that uses these functions and makes them async as well. Finally, it runs the unit tests again and saves the results
We evaluate PyMigTool using a new purely test-based evaluation setup, as opposed to the benchmark based evaluation we did in the first phase. This allows us to evaluate new migrations without the need for a ground truth developer migration. We evaluate PyMigTool on a set of 717 simulated migrations in 175 unique projects. By simulated migrations, we mean that this migration did not exist in the project’s history; instead, we find source libraries used in a project and migrate the project to an analogous target library. We ensure that the projects we choose have high test coverage, so that the evaluation is reliable.
We demonstrate twi different evaluation setups. First, we measure how well PyMigTool perform as it is on the simulated migrations, without any manual intervention. While the best case scenario is that PyMigTool can migrate the project without any issues, a more practical expectation is that PyMigTool would do most of the task correctly, but there are some parts that require manual intervention. Therefore, in a second evaluation, we measure how much manual work can fix the incorrect migrations. We randomly select a subset of 43 diverse migrations that were not fully correct, and we manually fix them.
We find that PyMigTool was able to successfully migrate 32% migrations without any manual intervention. We manually fix 43 of the remaining 484 migrations, and were able to fix 21 (49%) of them. In terms of modified lines of code, the manual work is only 14% of the total migration-related changes.
To summarize, the contributions of this paper are:
@@
    We conduct an initial detailed experiment to explore the use of LLMs for library migration on an existing migration benchmark that provides us a ground truth of previously performed developer migrations.
@@
    Based on the insights gained from our initial experiment, we build a CLI tool, PyMigTool, that automates the full library migration process using a combination of LLMs and post processing using static and dynamic program analysis.
@@
    We run an extensive test-based evaluation of PyMigTool on a set of 717 simulated migrations, and show that it works well with no or minimal manual editing for a diverse set of library pairs.
@@ 2.1 Library Migration
Library migration is the process of updating a software project to replace a used library with another one that provides similar functionality teyton2012mining . The source library is the one being replaced, and the target library is the one that replaces it teyton2012mining . One migration instance refers to a commit in a repository where a migration happened from a specific source library to a target library pymigbench , denoted using the notation source→target. For example, the commit b0607a26 in repository openstack/ironic is a retrying→tenacity migration.
A migration-related code change, or code change for brevity, is a minimal replacement of source library APIs with target library APIs that cannot be meaningfully reduced further without losing the semantics of the change pymigbench . A migration contains one or more code changes. The lines between the boxes in 1(c) show code changes. For example, segment P1 in 1(a) is replaced by segment D1 in 1(b).
We use subscripts pre, dev, and llm to denote data before migration, after developer’s migration, and after an LLM’s migration, respectively. For example, codepre, codedev, and codellm denote three states of a code. Changedev and Changellm denote code changes by developers and an LLM, respectively.
@@ 2.2 PyMigBench and PyMigTax
PyMigBench is a dataset of real-world Python library migrations pymigbench , mined from version control history. We use the latest available version, version 2.2.5 pymigtax in our experiments. The dataset has 321 migrations and 2,989 migration-related code changes. It includes a detailed description of each code change, including line numbers and API names, which facilitates our evaluation.
PyMigTax is a taxonomy of migration-related code changes pymigtax , built using the first version of PyMigBench. The authors validated its generalizability on additional third-party data. PyMigTax describes a code change based on three dimensions: (1) Program elements: the types of program elements involved in the change (e.g., function call and attribute); (2) Cardinality: how many source APIs are replaced with how many target APIs. For example, one-to-many cardinality means one source API is replaced with multiple target APIs; one-to-many, many-to-one, and many-to-many cardinalities are commonly referred to as higher-cardinality; and (3) Properties: Additional properties to describe the code change (e.g., element name change when the source and target APIs have different names.). The code change data in PyMigBench is annotated with the PyMigTax categories. We use PyMigTax categories (shown in Table 4) to understand the types of code changes LLMs can handle. The PyMigTax paper pymigtax has full category descriptions.
3 Phase one: Exploring the capabilities of LLMs for Python library migration
We use three off-the-shelf LLMs – LLama 3.1 llamapaper , GPT-4o mini openaimodels , and GPT-4o openaimodels – to perform migrations on the PyMigBench dataset. We refer to these models as LLama, Mini, and 4o in the rest of the paper, respectively. We evaluate the correctness of the migrations in two ways. First, we compare the LLM-migrated code with the developer-migrated code in PyMigBench and report correctness at two granularity levels: the full migration and the individual code changes. We use PyMigTax to identify which types of code changes the LLMs can/cannot handle. Second, for the subset of migrations with available unit tests, we run the available tests to assess the correctness of the migrated code.
@@ 2. RQ-A2
3.1 Experiment Setup
3.1.1 Models
Table 1: Description of the models we use in our experiment.
LLama 3.1 GPT-4o mini GPT-4o
Used version 3.1-70B-Instruct 2024-07-18 2024-08-06
Knowledge cutoff December 2023 October 2023 October 2023
Model Type Open-source Proprietary Proprietary
# training parameters 70 billion No official info No official info
Context length 128K tokens 128K tokens 128K tokens
Output limit 4096 tokens  [^1] LLama 3.1 has an input+output limit of 8192 tokens. Our prompt and response are approximately equal, therefore we consider the output limit to be half (4096). 16,384 tokens 16,384 tokens
Cost  [^2] Cost per million input token at the time of our experiments: July and August 2024 Free \$0.15 \$2.5
Table 1 shows the three off-the-shelf models we use for our library migration task. The first model is Meta’s LLama 3.1, a free and open-source LLM llamapaper . The other two models are proprietary: OpenAI GPT-4o mini and GPT-4o openaimodels . GPT-4o mini replaces OpenAI’s previous GPT-3.5 Turbo and is more cost-effective. GPT-4o is the most powerful of the OpenAI models, while also being the most expensive openaimodels . We use these models to migrate the code in 321 migrations from our PyMigBench dataset (pymigbench ).
3.1.2 Data preparation
At the time of conducting this experiment, two repositories from PyMigBench were no longer public, preventing us from using 2 migrations. We clone the remaining repositories containing 319 migrations. For each migration commit, we use PyDriller PyDriller to extract the content of each file that has recorded migration-related code changes, both before and after the migration. We use Python’s built-in ast.parse function docast to ensure that the files are syntactically valid code and discard 5 migrations having files with syntax errors. We conduct our experiments with the remaining 314 migrations containing 2,910 migration-related code changes.
3.1.3 Migration
For each file having code changes, we use the LLM prompt template in 1 to migrate the code. The prompt is designed to ensure that the LLM performs focused, controlled migrations while maintaining transparency. By asking the LLM to explain its changes, we ensure that the LLM justifies its modifications, which can be useful for a human evaluator reviewing the migration. The prompt also aims to restrict the LLM from making unrelated modifications such that we remain focused on the task of library migration.
Since LLMs are non-deterministic atil2024llmstabilitydetailedanalysis , we use a temperature of 0 and run each migration 10 times for each model. We calculate the difference between the runs to understand the variability in the generated code. We use git-diff gitdiff to compute a diff between each pair of the 10 runs and normalize the number of different lines by dividing it by the total number of lines in Filepre. We find that on average, two runs on the same file are only 6.6%, 4.0% and 4.8% different for LLama, Mini, and 4o, respectively. Given this low variability and the high effort involved in manual validation, we randomly select one run for each model to evaluate the LLMs’ results. Our artifact contains the LLM’s migrations for all 10 runs.
3.2 RQ-A1 How similar are the LLM migrations to the benchmark migrations?
3.2.1 RQ-A1 Approach
In this RQ, we assess each LLM’s migration correctness by comparing its code changes (Changellm) to those made by developers (Changedev), while manually judging potential alternative changes. A migration may involve multiple files, each with multiple migration-related code changes. We evaluate each code change individually, and aggregate the results to determine the correctness of the whole migration. We now explain the process using the example in 1(c) that shows a migration from the library requests to aiohttp.
@@ 3.2.1 RQ-A1 Approach
1The following Python code uses library <source-lib>.
2Migrate this code to use library <target-lib> instead.
3In the output, first explain the changes you made. Then, provide the modified code.
4
5Do not make any changes to the code that are not related to migrating between these two libraries.
6Do not refactor. Do not reformat. Do not optimize. Do not change coding style.
7
8Provided code:
9<premig-code>

### v2

@@ Abstract
Library migration is the process of replacing a library with a similar one in a software project. Manual library migration is time-consuming and error-prone, as it requires developers to understand the Application Programming Interfaces (API) of both libraries, map equivalent APIs, and perform the necessary code transformations. Due to the difficulty of the library migration process, most of the existing automated techniques and tooling stop at the API mapping stage or support a limited set of libraries and code transformations. In this paper, we develop an end-to-end solution that can automatically migrate code between any arbitrary pair of Python libraries that provide similar functionality. Given the promising capabilities of Large Language Models (LLMs) in code generation and transformation, we use them as the primary engine for migration. Before building the tool, we first study the capabilities of LLMs for library migration on a benchmark of 314 real-world library migrations. We find that the best performing LLM, GPT-4o, can correctly perform 57% of the migrations. We also identify post-processing steps that can further improve the performance. Based on this, we develop MigrateLib, a command-line application that combines the power of LLMs, static analysis, and dynamic analysis to provide accurate library migration. We evaluate MigrateLib on 717 migration scenarios in 175 real-world Python applications that are not from our benchmark. We find that MigrateLib can migrate 32% of the migrations with complete correctness without any developer intervention. Of the remaining migrations, 73% of the migration-related changes are automatically done by the tool for more than half of the projects.
@@ 1 Introduction
Library migration is the process of replacing a library with an alternative one in a software project. The process involves identifying the library usage in the codebase, finding the corresponding Application Programming Interfaces (APIs) in the new library, and replacing the old API usage with the new API usage. This can be tedious and error-prone, especially when the library is extensively used in a large codebase Kula et al. (2018).
@@
Researchers have proposed various techniques to automate the library migration process. Most of these techniques, however, are limited to only finding the analogous APIs (API mapping) Teyton et al. (2013); Alrubaye and Mkaouer (2018); Alrubaye et al. (2019b); Chen et al. (2019); Zhang et al. (2020); Di Rocco et al. (2025), but do not replace the old API usage in the codebase with the new one (code transformation). To find analogous APIs, some of these techniques mine existing migrations Teyton et al. (2013); Alrubaye and Mkaouer (2018); Alrubaye et al. (2019b) while others use machine learning models Chen et al. (2019); Di Rocco et al. (2025) and natural language processing techniques Ni et al. (2021). To the best of our knowledge, there are two techniques that actually perform code transformation Ni et al. (2021); Nikolov et al. (2025), but both techniques evaluate their approaches on a small set of selected library pairs without providing a tool that can be used in practice.
@@
In recent years, Large Language Models (LLMs) have shown that they can be effective in various software engineering tasks Ozkaya (2023); Fan et al. (2023); Wang et al. (2024), including the necessary upstream tasks for library migration, such as code generation Nguyen and Nadi (2022); Peng et al. (2023); Murali et al. (2023), code comprehension Nam et al. (2024), and code transformation Wei et al. (2023); Xia et al. (2023); Zhou et al. (2023a). This suggests that LLMs may be capable of performing library migration. However, studies have also identified limitations of LLMs in other software engineering tasks, especially generating incorrect or incomplete code Fan et al. (2023). These previous efforts suggest that LLMs may be capable of performing library migration and that LLM-based migration tooling could be effective. However, we still need further empirical evidence to support this claim.
In this paper, our goal is to investigate the effectiveness of LLMs in library migration to allow informed design decisions for LLM-based migration tooling. We choose Python as the programming language for our study, as it is one of the most popular programming languages and has a rich ecosystem of libraries. Accordingly, we approach this in three phases as illustrated in Figure 1. In Phase 1, our goal is to systematically investigate the capabilities and challenges of using LLMs for library migration. We conduct an experiment using 314 migrations from PyMigBench Islam et al. (2023), a Python library migration benchmark we previously built. We experiment with three LLMs (Llama 3.1, GPT-4o mini, and GPT-4o) and measure the correctness of the migrations by comparing the LLM’s migrations with the original developer’s migrations. We find that the three LLMs can correctly migrate 26%, 49%, and 57% of the migrations, respectively. Using an alternative test-based evaluation on a subset of the projects that have unit tests, we find that 36%, 52%, and 64% of the LLM migrations maintain the same test results as the original developer migrations for the three LLMs, respectively. We also study the types of migration-related code changes LLMs excel at vs. struggle with. Overall, we find that LLMs are capable of library migration, and as expected, the more powerful LLM (GPT-4o) performs best. We also identify four major challenges that make solely relying on LLMs in a practical tool less effective: (1) The files that need migration are not always known beforehand. (2) All LLMs often return only part of the code, skipping parts of the code that do not require migration. (3) For a given project, a file that we migrate may have other files that depend on it which may require changes during migration. Specifically, if any functions are made asynchronous during the migration, all parts of the code that use those functions must also be made asynchronous. (4) The inconsistencies between the versions between the libraries used in the project and those assumed by the LLM can lead to incorrect migrations.
@@
We find that the identified limitations from Phase 1 can be addressed by improving the prompt, some pre-processing and post-processing the LLM results. Accordingly, in Phase 2, we build a CLI tool, MigrateLib, that automates the full library migration process using LLMs and post-processing using static and dynamic program analysis. MigrateLib takes in a complete Python project with an existing test suite and automatically performs the migration for the specified library pair. By building a dynamic call graph, MigrateLib first identifies all files that require migration and sends those to the LLM via a prompt that instructs the LLM to perform the migration. MigrateLib then verifies the results of the migration by comparing the test results before and after migration. If any previously passing tests now fail, MigrateLib performs a post-processing step where it checks if the LLM skipped any code and merges it back. It also detects whether any functions were made async during the migration, and if so, finds other parts of the project that use those functions and makes them async as well. Finally, it runs the unit tests again to determine the final migration status.
PyMigBenchPhase 1 Explore LLMs for Library MigrationLlama 3.1GPT-4o miniGPT-4oPhase 2 Build MigrateLib Phase 3 Evaluate MigrateLibGPT-4oSimulated migrations314 migrations717 migrations
Figure 1: Overview of the contributions.
In Phase 3, we evaluate MigrateLib using a new purely test-based evaluation setup, as opposed to the benchmark-based evaluation we did in Phase 1. This allows us to conduct a larger-scale evaluation and also evaluate new migrations without the need for ground truth developer changes. We run MigrateLib on a set of 717 simulated migrations in 175 unique projects. We define a simulated migration as one that did not exist in the project’s history. Specifically, we find source libraries used in a project and migrate the project to an analogous target library. We ensure that the projects we choose have high code coverage (above 80%), so that the evaluation is reliable. Given that GPT-4o had the best performance in Phase 1, we conduct this evaluation only using GPT-4o. We find that MigrateLib is able to successfully complete 32% migrations without any manual intervention (i.e., all tests that pass before migration still pass after migration). To understand how much manual work is required to fix the remaining 484 migrations that the LLM left in an incomplete stage, we randomly select a subset of 43 diverse migrations and attempt to manually fix them. We measure the amount of manual work required to complete these migrations using the number of lines of code we need to manually edit. We find that we are able to fix 24 (56%) of our sampled migrations, where the manual work is only 27% of the total changes needed for the migration. The main reason we could not fix some of the remaining migrations is that the projects had complex logic that required a deep understanding of the project context, which was difficult to achieve.
Overall, our work provides a combination of empirical and engineering contributions. This paper is an extended version of our prior conference paper Islam et al. (2025) that only focused on Phase 1. Phases 2 and 3 are novel contributions of this work. To summarize, the contributions of this paper are:
@@
    An empirical study to explore the use of LLMs for library migration by comparing LLM changes to 314 benchmark developer changes. Our analysis systematically highlights the types of migration-related code changes LLMs excel at vs. struggle with.
@@
    We design and build an open-source end-to-end CLI tool, MigrateLib, that automates the full library migration process using a combination of LLMs and post-processing using static and dynamic program analysis.
@@
    We design a new evaluation setup that allows us to evaluate MigrateLib on a large scale and also on newer projects. We run an extensive test-based evaluation of MigrateLib on a set of 717 simulated migrations, and show that it works well with no or minimal manual editing for a diverse set of library pairs.
The source code of MigrateLib, the experimental data, and the detailed results of our experiments are available at https://figshare.com/account/articles/29602517.
@@ 2.1 Library migration
Library migration is the process of updating a software project to replace a used library with another one that provides similar functionality Teyton et al. (2012). The source library is the one being replaced, and the target library is the one that replaces it Teyton et al. (2012). One migration instance refers to a commit in a repository where a migration happened from a specific source library to a target library Islam et al. (2023), denoted using the notation source→target. For example, the commit b0607a26 in repository openstack/ironic is a retrying→tenacity migration.
A migration-related code change, or code change for brevity, is a minimal replacement of source library APIs with target library APIs that cannot be meaningfully reduced further without losing the semantics of the change Islam et al. (2023). A migration contains one or more code changes. The lines between the boxes in 2(c) show code changes. For example, segment P1 in 2(a) is replaced by segment D1 in 2(b).
We use subscripts pre, dev, and llm to denote data before migration, after developer’s migration, and after an LLM’s migration, respectively. For example, Codepre, Codedev, and Codellm denote three states of a piece of code. Changedev and Changellm denote code changes by developers and an LLM, respectively.
@@ 2.2 PyMigBench and PyMigTax
We use two of our prior artifacts: PyMigBench Islam et al. (2023) and PyMigTax Islam et al. (2024), in this work which we briefly describe next. PyMigBench is a dataset of real-world Python library migrations, which we mined from open source GitHub repositories. We use the latest version, version 2.2.5, in our experiments in this paper. The dataset has 321 migrations and 2,989 migration-related code changes. It includes a detailed description of each code change, including line numbers and API names, which facilitates our evaluation.
PyMigTax is a taxonomy of migration-related code changes Islam et al. (2024), built using the first version of PyMigBench. In our prior work, we validated its generalizability on additional third-party data. PyMigTax describes a code change based on three dimensions: (1) Program elements: the types of program elements involved in the change (e.g., function call and attribute); (2) Cardinality: how many source APIs are replaced with how many target APIs. For example, one-to-many cardinality means one source API is replaced with multiple target APIs; one-to-many, many-to-one, and many-to-many cardinalities are commonly referred to as higher-cardinality; and (3) Properties: Additional properties to describe the code change (e.g., element name change when the source and target APIs have different names.). The code change data in PyMigBench is already annotated with the PyMigTax categories. We use PyMigTax categories (shown in Table 3) to understand the types of code changes LLMs can handle. The PyMigTax paper Islam et al. (2024) has full category descriptions.
3 Phase 1: Exploring LLMs for library migration
We first systematically explore the capabilities of LLMs for library migration, while understanding how they handle different types of migration-related code changes. We first explain our setup and methods, and then present the results.
3.1 Experiment setup
@@
We start with the 321 migrations in PyMigBench for our experiments, then discard 2 migrations whose repositories are no longer public, and 5 migrations where either the pre-migration or post-migration file has syntax errors. We conduct our experiments with the remaining 314 migrations containing 2,910 migration-related code changes.
@@
The following Python code uses library <source-lib>. Migrate this code to use library <target-lib> instead. In the output, first explain the changes you made. Then, provide the modified code.
Do not make any changes to the code that are not related to migrating between these two libraries. Do not refactor. Do not reformat. Do not optimize. Do not
change coding style.
Provided code:
<premig-code>
3.1.1 Performing migration
We use three off-the-shelf LLMs – Llama 3.1-70B-Instruct Dubey et al. (2024), GPT-4o-mini-2024-07-18, and GPT-4o-2024-08-06 OpenAI (2024) – to perform migrations on the PyMigBench dataset. We refer to these models as Llama, Mini, and 4o in the rest of the paper, respectively.
For each file having code changes in a given migration, we use the LLM prompt template in 1 to migrate the code. The prompt is designed to ensure that the LLM performs focused, controlled migrations while maintaining transparency. By asking the LLM to explain its changes, we ensure that the LLM justifies its modifications, which can be useful for a human evaluator reviewing the migration. The prompt also aims to restrict the LLM from making unrelated modifications such that we remain focused on the task of library migration.
Since LLMs are non-deterministic Atil et al. (2024), we use a temperature of 0 and run each migration 10 times for each model. To understand the extent of variability in the generated code, we calculate the difference between the runs. We use git-diff to compute a diff between each pair of the 10 runs and normalize the number of different lines by dividing it by the total number of lines in Filepre. We find that on average, two runs on the same file are only 6.6%, 4.0% and 4.8% different for Llama, Mini, and 4o, respectively. Given this low variability and the high effort involved in manual validation, we randomly select one run for each model to evaluate the LLMs’ results. Our artifact contains all LLM migrations for all 10 runs.
3.1.2 Matching LLM migrations with developer migration
In RQ-1.1, we compare the LLM migrations with the developer migrations in PyMigBench. We do this for each code change, and then aggregate the results at the migration level, while also condidering the additional non-migration-related changes made by the developer and the LLM. We now explain the process using the example in 2(c) that shows a migration from the library requests to aiohttp.